# AutoElicit: Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v4

@@ A.7.4 Extracting the in-context model’s hidden prior and posterior
We can then measure, as in Section 5.5, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.5) between this sample and the extracted in-context sample and through visual inspection. By doing this, we test whether the language model is using the same distribution we elicit from it as the one it uses in-context. The results of this are presented in Table 1.
@@
A.7.5 Calculating Bayes factor
@@
For a single dataset, we start by randomly selecting $25$ data points, $5$ times, and consider the prior predictive distribution. For prior elicitation this involves sampling the prior predictive distribution (drawing $500$ samples for each split), and for in-context learning this involves prompting a language model for a prediction for each of the $100$ task descriptions (producing $100$ samples for each of the splits).
@@
We then report the mean and standard deviation of the Bayes factor over the splits in Figure 7, which helps us to decide, given a dataset, which method is the most appropriate; assuming that in-context learning approximates Bayesian inference.
@@ A.7.6 Given a new dataset
In all of our experiments, we asked the language model to mimic a linear model when predicting in-context, and to provide priors for a linear model when performing prior elicitation.
@@
Now, to measure whether prior elicitation or in-context learning is better for a given task (assuming the language model is approximating Bayesian inference):
@@
    Then, to perform in-context learning, use a similar task description to above but with a different system and user role that specifies the language model should make predictions instead of providing prior distributions. With this new set of descriptions, again ask a language model to rephrase them.
@@
To empirically check whether the language model is approximating Bayesian inference to produce its in-context predictions or to test whether it is being faithful to the priors we elicit, we can follow the below procedure:
@@
For this, we can follow the below procedure:
@@ A.8 Dataset memorisation
We then use the Levenshtein distance (Levenshtein, 1965; 1966), a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00, suggesting that the strings are equal, and a maximum value of the longest length of the two strings.
@@
Table 2: GPT-3.5-turbo header completion test results. Normalised Levenshtein distance between the correct header completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
@@
The values in Table 2 demonstrate that the language model has completely memorised the first few rows of the Breast Cancer and Heart Disease datasets, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For the Diabetes and Hypothyroid datasets, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the completion from the language model. The UTI task has not been tested, as it is a private dataset.
Table 3: GPT-3.5-turbo row completion test results. Normalised Levenshtein distance between the correct row completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
@@
We supplement these results with those of Table 3, which shows the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 3 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases.
@@ A.8.1 How does this affect our results?
Within our experiments, we normalise the training and testing data, as well as sample random observations from the dataset. Because of this, observations predicted during in-context learning will likely not have been memorised, since they are normalised based on a training data subset and are selected from random rows of the dataset. As shown, although the language model can complete the first few rows of the datasets with some success, it is unable to complete random rows (Table 3). Therefore, we hypothesise that any memorisation of the tested datasets has little impact on the in-context predictions.
Further, since for prior elicitation we only prompt the language model for a prior distribution of a predictive model given the feature names and a description of the dataset, the memorisation of rows from the data is unlikely to have had an impact on these results. In addition, we see that elicited priors provide improved predictions over the uninformative prior for the private and synthetic datasets, as well as the public datasets.
@@ A.9 Additional AutoElicit results
In Figure 9 we present the accuracy distributions from the experiment in Figure 2, from Section 5.2.
@@
For Hypothyroid prediction, DeepSeek-R1-32B-Q4 provides the best accuracy by a large margin. Of the OpenAI models, smaller training sizes favour GPT-4-turbo, whilst larger training sizes favour, GPT-3.5-turbo. This is because the GPT-4-turbo-elicited prior was too strict to allow for updated posterior parameters, but was more informative than an uninformative prior for smaller training sizes.
@@
A.9.1 Priors elicited from the synthetic data
@@
A.9.2 UTI accuracy with the label collection context
@@
In Figure 11 we plot the time it took to collect the given number of labels tested in the experiments in Section 5.2 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve a greater accuracy with fewer data points, and therefore less data collection time.
@@
It is important to note that the accuracy in Figure 11 is not achieved as a result of training on the earliest collected data points. This is because in our analysis we perform cross-validation and use a subset of the data from a later time period (Section A.7.1). Changes in the in-home devices were made during the study, such as the addition of a device to collect sleep information, which we use within the dataset evaluated in the main results. However, if all data collection devices had been used from the beginning of the study, we could expect the accuracy in time to resemble the figure presented here.
@@ A.11 Task descriptions with more information
Figure 12: Mean squared error using the GPT-3.5-turbo-elicited priors for the task: $y=2x_{1}-x_{2}+x_{3}$ where we have used increasingly more detailed task descriptions. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$. Within each task description we provide the language model with increasingly more information about the task, as to test how the task description effects the priors elicited. Complete descriptions of the task descriptions given here are provided in Section A.11. The lines depict the average mean posterior mean squared error (MSE) over the 5 dataset splits, whilst the error bars represent the 95% confidence interval.
To understand how the detail of the task description affects the performance of the priors elicited from the language model, we studied the synthetic task $y=2x_{1}-x_{2}+x_{3}$, and describe the relationship between features and targets in increasing detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 12, we included the following text in the task description that was provided to the language model:
@@
This demonstrates the sometimes unexpected behaviour of language models. Here, the LLM produced Gaussians for the prior distribution with small standard deviations (reflecting a high confidence in the distribution) and means that imply the language model knew the underlying equation defining the relationship between the target and features. We might have expected that after only being told the relationship was linear, the language model would produce Gaussians with 00 mean and standard deviation $1$, reflecting the uncertainty in the exact relationship between each feature and the target.
Interestingly, although “Three feature relationships” and “Full equation” contain the same information written differently, the language model provided better prior distributions when given the full equation. Fully specifying the equation is not realistic in a real-world setting where we are not sure about the exact relationship between features and targets, but this provides an example of how we can make the task description as information dense as possible.
Figure 13: GPT-3.5-turbo-elicited parameter distribution as increasingly more information is included. These histograms show the distribution of parameter values for the features as more expert information is provided. In each row, we provide information in the task description of an additional feature, with the green arrow indicating the effect this has on the parameter distribution. In particular, in the first row we provide information about the relationship between the target and the first feature, in the second row we provide information about the second feature, and the in the third row we provide information about the third feature.
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 13 we show the parameter distributions elicited from the language model for the different descriptions we provide. As we provide additional information on each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions. This experiment provides a controlled setting that supplements the results in Section 5.3.
@@ A.12 How many ways should we describe a task?
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher-dimensional dataset.
This figure shows that as the number of task descriptions increases, we see smaller changes in the energy distance and reduced variability in the samples. This suggests that increasing the number of task descriptions provides diminishing returns.
@@
A.13 Is the language model being consistent during in-context learning?
@@
Figure 15 visualises the different prior parameter distributions for all of the tested datasets in each feature dimension. Notable are the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 5.5. In the in-context case, prior distributions often have modes close to 00, suggesting that the internal model is not sure about the relationship between features and targets. The elicited priors often represent a stronger belief in the correlation between features and targets, with most histograms achieving modes greater than or less than 00.
@@
A.14 Empirically, is the language model approximating Bayesian inference?

### v5

@@ A.7.4 Extracting the In-context Model’s Hidden Prior and Posterior
We can then measure, as in Section 5.5, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.5) between this sample and the extracted in-context sample, and through visual inspection. By doing this, we test whether the language model is using the same distribution we elicit from it as the one it uses in-context. The results of this are presented in Table 1.
@@
A.7.5 Calculating the Bayes factor
@@
For a single dataset, we start by randomly selecting $25$ data points, $5$ times, and consider the prior predictive distribution. For prior elicitation, this involves sampling the prior predictive distribution (drawing $500$ samples for each split), and for in-context learning, this involves prompting a language model for a prediction for each of the $100$ task descriptions (producing $100$ samples for each of the splits).
@@
We then report the mean and standard deviation of the Bayes factor over the splits in Figure 7, which helps us to decide which method is most appropriate for a dataset; assuming that in-context learning approximates Bayesian inference.
@@ A.7.6 Given a New Dataset
To measure whether prior elicitation or in-context learning is better for a given task (assuming the language model is approximating Bayesian inference):
@@
    Then, to perform in-context learning, use a similar task description to the above but with a different system and user role that specifies the language model should make predictions instead of providing prior distributions. With this new set of descriptions, again ask a language model to rephrase them.
@@
To empirically check whether the language model is approximating Bayesian inference to produce its in-context predictions or to test whether it is being faithful to the priors we elicit, we can follow the procedure below:
@@
For this, we can follow the procedure below:
@@ A.8 Dataset Memorisation
We then use the Levenshtein distance (Levenshtein, 1965, 1966), a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00, suggesting that the strings are equal, and a maximum value of the longest length of the two strings.
@@
Table 2: GPT-3.5-turbo header completion test results. Normalised Levenshtein distance between the correct header completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal, and a value of $1$ means that they were different at all positions.
@@
The values in Table 2 show the Levenshtein distance between the true dataset header completion and the one made by the LLM. This demonstrates that the language model has completely memorised the first few rows of the Breast Cancer and Heart Disease datasets, shown by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For the Diabetes and Hypothyroid datasets, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the completion from the language model. The UTI task has not been tested, as it is a private dataset.
Table 3: GPT-3.5-turbo row completion test results. Normalised Levenshtein distance between the correct row completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal, and a value of $1$ means that they were different at all positions.
@@
We supplement these results with those of Table 3, which shows the results of the row completion test. Here, we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 3 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases.
@@ A.8.1 How Does this Affect our Results?
Within our experiments, we normalise the training and testing data and sample random observations from the dataset. Because of this, observations predicted during in-context learning will likely not have been memorised, as they are normalised based on a training data subset and are selected from random rows of the dataset. Also, although the LLM can complete the first few rows of the datasets with some success, it is unable to complete random rows. Therefore, we hypothesise that any memorisation of the tested datasets has little impact on the in-context predictions.
Further, since for prior elicitation we only prompt the language model for a prior distribution of a predictive model given the feature names and a description of the dataset, the memorisation of rows from the data is unlikely to have had an impact on these results. In addition, we see that elicited priors provide improved predictions over the uninformative prior for the private and synthetic datasets (with no dataset memorisation), as well as the public datasets.
@@ A.9 Additional AutoElicit Results
A.9.1 Accuracy Distributions
In Figure 9, we present the accuracy distributions from the experiment in Figure 2, from Section 5.2.
@@ A.9.1 Accuracy Distributions
For Hypothyroid prediction, DeepSeek-R1-32B-Q4 provides the best accuracy by a large margin. Of the OpenAI models, smaller training sizes favour GPT-4-turbo, whilst larger training sizes favour GPT-3.5-turbo. This is because the GPT-4-turbo-elicited prior was too strict to allow for updated posterior parameters, but was more informative than an uninformative prior for smaller training sizes.
@@
A.9.2 Priors Elicited from the Synthetic Data
@@
A.9.3 UTI Accuracy with the Label Collection Context
@@
In Figure 11, we plot the time it took to collect the given number of labels tested in the experiments in Section 5.2 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve greater accuracy with fewer data points, and therefore less data collection time.
@@
It is important to note that the accuracy in Figure 11 is not achieved as a result of training on the earliest collected data points. This is because in our analysis, we perform cross-validation and use a subset of the data from a later time period (Section A.7.1). Changes in the in-home devices were made during the study, such as the addition of a device to collect sleep information, which we use within the dataset evaluated in the main results. However, if all data collection devices had been used from the beginning of the study, we could expect the accuracy in time to resemble the figure presented here.
A.9.4 Another Uninformative Prior
For additional context to the results provided in Figure 2, we present the posterior performance of a different uninformative prior. In this further case, we sample $100$ values from a standard normal distribution ($\mu_{k}\sim\mathcal{N}(\mu|0,1)$), and we use these as means of $100$ Gaussian distributions ($\theta_{k}\sim\mathcal{N}(\theta|\mu_{k},1)$). With these distributions, we construct a mixture prior. As we keep the standard deviation at 1 for each component, the resulting uninformative prior has a greater range in the parameter space.
Figure 12 shows the results of this experiment compared to those achieved using AutoElicit. This new uninformative prior leads to no noticeable difference in the results for the synthetic task, Diabetes task, and Breast Cancer task. However, we see marginally better accuracy on the UTI task than the previous uninformative prior, and on the Heart Disease and Hypothyroid tasks, we see greater improvements. Despite this, DeepSeek-R1-32B-Q4 or GPT-3.5-turbo continue to provide more informative priors.
This suggests that the performance improvement we see from using AutoElicit over an uninformative prior is not due to its complexity.
Figure 12: Comparing AutoElicit to a mixture of uninformative priors. This figure presents a linear model’s posterior accuracy or mean squared error after observing the given number of training data points, over $10$ repeats. Here, we compare the posterior performance of using AutoElicit with two LLMs to that of a mixture of uninformative priors, where the means of each component are sampled from a standard Gaussian distribution, as well as the uninformative prior presented in Figure 2.
Figure 13: AutoElicit and baseline from Gouk & Gao (2024). This figure demonstrates a linear model’s posterior accuracy or mean squared error after observing the given number of training data points over 10 repeats. The solid lines correspond to the results from AutoElicit, whilst the dashed lines correspond to the results from Gouk & Gao (2024). The left-most two datasets are unseen by an LLM, whilst the right-most four datasets are publicly available and likely have samples within the LLM’s training data. In these cases, the LLM could be reproducing real data from the task, artificially improving the results for the method proposed by Gouk & Gao (2024).
A.9.5 Comparing with Gouk & Gao (2024)
The method presented by Gouk & Gao (2024) generates synthetic data for a predictive task using an LLM. Using this data, they provide an additional likelihood term when training linear classification models. Specifically, they use an LLM to generate feature values based on a task description, which are labelled using zero-shot in-context learning. The samples generated are provided alongside real data during the training of a linear predictive model. However, this means that this method is susceptible to LLM data memorisation, putting it at risk of artificially providing improved posterior results on public datasets. Given this is a concern discussed in Bordt et al. (2024) and Appendix A.8, during our evaluation, we split the results by those achieved on publicly available tasks and those on privately available tasks. Additionally, to fully compare this method with AutoElicit, we extended their work to regression tasks. Figure 13 shows the results of this experiment.
In the public tasks, we see that the performance of the approach presented in Gouk & Gao (2024) varies considerably by LLM and task. It underperforms compared to the uninformative prior on five occasions. On the two occasions it outperforms AutoElicit (ours), it achieves surprisingly good accuracy or mean squared error (Heart Disease with GPT-4-turbo and Diabetes with GPT-3.5-turbo), suggesting that it might be reproducing parts of the public datasets rather than generating new samples. Given our analysis in Appendix A.8, showing the memorisation ability of LLMs, and the results of the method given in Gouk & Gao (2024) on private data, we suspect the performance of this approach is artificially improved by reproducing the true dataset. Since for AutoElicit, we prompt the language model for a prior distribution over the parameters of a predictive model, using only the feature names and a description of the dataset, the memorisation of data points is unlikely to have had an impact on our method’s results (Appendix A.8.1).
Moreover, in both of the private datasets (the two left-most plots), where there is no risk of LLM memorisation, AutoElicit results in improved posterior accuracy and mean squared error. On the synthetic task, our improvement is orders of magnitude better, and on the UTI task, our method significantly increases accuracy at all training sizes. In both cases, the approach proposed by Gouk & Gao (2024) underperforms compared to an uninformative prior.
@@ A.11 Task Descriptions with More Information
Figure 14: Mean squared error using the GPT-3.5-turbo-elicited priors for the task: $y=2x_{1}-x_{2}+x_{3}$ where we have used increasingly more detailed task descriptions. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$. Within each task description, we provide the language model with increasingly more information about the task to test how the task description affects the priors elicited. Complete descriptions of the task descriptions given here are provided in Section A.11. The lines depict the average mean posterior mean squared error (MSE) over the 5 dataset splits, whilst the error bars represent the 95% confidence interval.
Figure 15: GPT-3.5-turbo-elicited parameter distribution as increasingly more information is included. These histograms show the distribution of parameter values for the features as more expert information is provided. In each row, we provide information in the task description of an additional feature, with the green arrow indicating the effect this has on the parameter distribution. In particular, in the first row we provide information about the relationship between the target and the first feature, in the second row we provide information about the second feature, and in the third row we provide information about the third feature.
Figure 17: AutoElicit with different numbers of mixture components in the prior. This figure demonstrates a linear model’s posterior accuracy or mean squared error after observing the given number of training data points, over $10$ repeats. The rows of the figure correspond to the different language models tested, whilst the colour of the lines indicates the number of mixture components used in the prior. Using $16$ mixture components corresponds to using a product of $4$ system roles and $4$ user roles when eliciting priors.
To understand how the detail of the task description affects the performance of the priors elicited from the language model, we studied the synthetic task $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in increasing detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 14, we included the following text in the task description that was provided to the language model:
@@
This demonstrates the sometimes unexpected behaviour of language models. Here, the LLM produced Gaussians for the prior distribution with small standard deviations (reflecting a high confidence in the distribution) and means that the language model knew the underlying equation defining the relationship between the target and features. We might have expected that after only being told the relationship was linear, the language model would produce Gaussians with 00 mean and standard deviation $1$, reflecting the uncertainty in the exact relationship between each feature and the target.
Interestingly, although “Three feature relationships” and “Full equation” contain the same information written differently, the language model provided better prior distributions when given the full equation. Fully specifying the equation is not realistic in a real-world setting where we are unsure about the exact relationship between features and targets, but this provides an example of how we can make the task description as information-dense as possible.
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 15, we show the parameter distributions elicited from the language model for the different descriptions we provide. As we provide additional information on each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. Since the language model knows nothing about this synthetic task beforehand, this experiment demonstrates the effectiveness of using the AutoElicit framework to convert expert information as natural language into prior probability distributions. This experiment provides a controlled setting that supplements the results in Section 5.3.
@@ A.12 How Many Ways Should We Describe a Task?
It is interesting that Breast Cancer benefited the most from increasing the number of task descriptions, since this dataset has the largest number of features. It suggests that more task descriptions are required for datasets with high-dimensional data. This makes intuitive sense since we might expect a larger mixture of Gaussian components to better describe the prior over for a higher-dimensional dataset.
@@
To supplement this, we calculate the posterior performance with varied numbers of mixture components in our prior, shown in Figure 17. We see little variation (except for Breast Cancer with DeepSeek-R1-32B-Q4) in the accuracy or mean squared error as we increase the number of components in the prior, suggesting our framework is robust to the number of task descriptions. We hypothesise that a greater number of mixture components makes the framework more resilient to LLMs that frequently hallucinate, especially on tasks with large numbers of features (such as Breast Cancer).
Figure 16 showed that the density of the priors changed with the number of mixture components, however, this has a small effect on the posterior performance. This is likely because the sample space that provides good posterior performance is covered by the majority of elicited components, whilst the large changes in Figure 16 are due to more unique components.
@@
A.13 Are AutoElicit Priors Interpretable?
As the density of the distributions in Figure 18 are used for the prior directly, they allow for interpretable predictions. Considering the AutoElicit priors, for linear models, each feature’s prior corresponds to its direction and importance when predicting the target. On the synthetic task (row 1), the prior distributions concentrate on the parameters that we expect. For the UTI task (row 2), the LLM provided prior distributions that suggest Feature 2 was negatively predictive of a positive UTI, whilst Feature 10 is a strong positive predictor. For Breast Cancer (row 6), the LLM produced two modes in its Feature 2 prior. Here, the LLM was either unsure of its belief, and so the elicited prior changed between task descriptions, or the LLM was unsure of the feature’s direction of correlation but guessed that the feature was important.
A chain-of-thought LLM such as DeepSeek-R1-32B-Q4 could help us better understand the elicitation process.
A.14 Is the Language Model Being Consistent During In-context Learning?
@@
Figure 18 visualises the different prior parameter distributions for all of the tested datasets in each feature dimension. Notable are the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 5.5. In the in-context case, the prior distributions often have modes close to 00, suggesting that the internal model is not sure about the relationship between features and targets. In contrast, the elicited priors often represent a stronger belief in the correlation between features and targets, with most histograms achieving modes greater than or less than 00.
@@
A.15 Empirically, Is the Language Model Approximating Bayesian Inference?